### v1

# Can Large Language Models Address Open-Target Stance Detection?

Abstract
Stance detection (SD) assesses a text’s position towards a target, typically labeled as "favor," "against," or "neutral." We introduce Open-Target Stance Detection (OTSD), where targets are neither seen during training nor provided as input. Evaluating Large Language Models (LLMs) like GPT-3.5, Llama 3, and Mistral, we compare their performance with the Target-Stance Extraction (TSE) approach, which has the advantage of using predefined targets. LLMs perform better than TSE in target generation when the real target is explicitly and not explicitly mentioned in the text. For stance detection, LLMs perform better in explicit scenarios but fail in non-explicit ones. [^1] We will release the code and the dataset.
Can Large Language Models Address Open-Target Stance Detection?
Abu Ubaida Akash  and Ahmed Fahmy  and Amine Trabelsi {akaa2803, mosa2801, amine.trabelsi}@usherbrooke.ca Université de Sherbrooke
1 Introduction
Stance detection (SD) is the task of determining the position of a text or person towards a certain target, often split into “favor” and “against” with an optional third split being “neutral”, “none”, or “neither” (Sutter et al., 2024). The target word can be mentioned in the text explicitly, or sometimes the idea of the target word can be conveyed indirectly (Küçük and Can, 2020). In Zero-shot Stance Detection (ZSSD), the model predicts the stance for targets that it has not seen during the training phase. This task is essential because collecting training data for every possible topic or target is impractical (Allaway et al., 2021). While recent research has focused on ZSSD (Zhang et al., 2023; Li et al., 2023b; Wen and Hauptmann, 2023; Liang et al., 2022a; Zhu et al., 2022; Luo et al., 2022; Xu et al., 2022; Liang et al., 2022b; Li and Yuan, 2022), most studies assume that the target is known or manually identified and given as input, a situation that is uncommon in real-world applications. In practice, the target is unknown, and it is not necessarily mentioned explicitly in the text. Additionally, target annotation is a tedious and expensive task in SD (Küçük and Can, 2020; Sutter et al., 2024).
In this paper, we focus on a different yet challenging task we refer to as Open-Target Stance Detection (OTSD). In this task, the target is neither seen during training (zero-shot) nor provided as input to the model. In OTSD, one important challenge lies in identifying the target from the text rather than having it provided as input. The target may be mentioned explicitly or not. The stance is predicted with respect to the produced target.
There has been one notable attempt to partially address this real-world setting of SD. Li et al. (2023a) generate targets from the input text and later detect stance based on these targets. However, before starting the SD task, they map these generated targets to a predefined list of golden targets (details in §2) to ensure exact wording and facilitate evaluation. They called their approach Target-Stance Extraction in the Wild (TSE). This approach does not fully align with our OTSD setting, which assumes no input target information is given during the whole process, and is impractical due to the need for a comprehensive list of all possible targets. Moreover, our experiment reveals a performance gap in Li et al. (2023a)’s approach when used on text with explicit or non-explicit target mentions. This aspect was not highlighted by the authors and warrants further investigation in the context of OTSD.
In this work, we present and explore the OTSD task, focusing on its two main steps: Target Generation and Stance Detection (see Eq. 2). As a zero-shot task for both target and stance, Large Language Models (LLMs) appear well-suited for this challenge. Therefore, we examine the performance of LLMs and compare it to TSE, the primary existing work by Li et al. (2023a), which benefits from using a predefined list of targets during the process (see Eq. 1). Our empirical study aims to address the following research questions: [RQ#1] How do proprietary and open LLMs—specifically GPT-3.5 (Brown et al., 2020), Llama 3 (Touvron et al., 2023), and Mistral (Jiang et al., 2023)—perform in open target generation compared to the TSE approach when the real target is explicitly or not explicitly mentioned in the text for stance classification, and how do they compare to each other? [RQ#2] How do the same LLMs perform compared to TSE in stance detection (in the context of OTSD) for both explicit and non-explicit cases, and how do they compare to each other?
Our contributions consist of introducing the task of OTSD and conducting experiments and answer these research questions.
2 Target-Stance Extraction (TSE)
In TSE, Li et al. (2023a) demonstrate target extraction and stance detection tasks in both in-target and zero-shot settings. For our comparison, we only consider their zero-shot setting. In TSE zero-shot, they first generate the target using keyphrase generation models, map the generated targets (multiple targets per sample) to a predefined list of golden targets to find the most similar one, and finally use it to detect the stance of the text toward the target (Eq. 1). The best match from their keyphrase model’s output, marked as TSE-BestGen, is used for our comparison. The best mapped target (TSE-mapped) is considered the predicted target. For example, given the text "Embracing different faiths teaches us empathy and respect. Let’s celebrate our diverse beliefs and foster unity. #Faith #Unity #Respect", TSE primarily generates targets such as Peace, Religious diversity, Respect, etc.. From a predefined list of possible targets (e.g., Face Mask, Atheism, Donald Trump, etc.), Atheism matches most closely with Religious diversity and is considered as the final predicted target. In OTSD, we aim at directly generating Religious diversity or a very semantically similar phrasing, avoiding the constraints of a predefined list. In the second step, they detect the stance toward their generated target.
3 Open-Target Stance Detection (OTSD)
3.1 Task Definition
Given an input text $x$, previous work (Eq. 1) generates a target $t^{\prime}$ and maps it to a predefined list of targets $k$, resulting in a mapped target $t$. Finally, they detect the stance $y$ given $x$ and $t$. OTSD task (Eq. 2) is to detect the stance $y$ from the input text $x$ and the generated target $t^{\prime}$, eliminating the need for a predefined list $k$.
[…]
3.2 Approach
To address the research questions outlined in §1, our approach to OTSD leverages the zero-shot learning capabilities of LLMs. We adopt the "Task Definition" prompting strategy for the OTSD task, as it has been shown by Cruickshank and Ng (2023) to outperform other strategies—such as "Task-only", "Context Analyze", "Context Question", "Zero-shot CoT", and "CoDA"—particularly for smaller models (7B-8B parameters) across most datasets. Given that this is an open-target task, we aim to use a straightforward prompting approach to maintain fairness and generalizability; we do not tailor the prompt for specific targets or include demonstrations or in-context examples, as this remains a zero-shot task.
Since LLMs are sensitive to the length of generated text, we limit the target length to ensure a fair comparison with TSE in terms of target generation quality. To determine the optimal length, we test various target length instructions across all LLMs (§4.1) on 100 randomly selected samples from our dataset (§4.1), aiming to achieve an average length and standard deviation that closely match those of the golden targets. Based on our empirical study, we set the maximum target length to 5 for GPT-3.5 and 4 for Llama and Mistral to obtain an average length and standard deviations close to those of gold targets (explicit: $3.78\pm 1.5$; non-explicit 2.66±1.76)2.66$\pm$ 1.76)2.66 ± 1.76 ). Finally, we employ two different prompting approaches for target generation (TG) and stance detection (SD).
TG+SD (Two-Step Approach)
In this approach, we first generate the target and then sequentially detect the stance of the text toward the generated target. By directing the LLM to focus on one task at a time, this method may reduce the cognitive load on the model.
TG&SD (Single-Step Joint Approach)
This approach combines target generation and stance detection into a single step using a unified prompt. By jointly generating the target and stance, the model can develop a more holistic understanding of the relationship between the text, target, and stance.
The prompts and their design justifications can be found in Appendix A.
4 Experiments
In the OTSD task, we compare our approaches with TSE in both TG and SD. To ensure sensitivity and robustness, we run our experiments three times independently and report the average results  [^2] Score variations are negligible, up to four decimal places.. Since OTSD is a generative task, we primarily compare our results with TSE-BestGen, despite its reliance on a predefined list of golden targets. Additionally, we consider TSE-mapped as the current best score, regardless of task type. Details about our hardware settings can be found in Appendix D.
4.1 Dataset and Model
We use the same dataset as TSE’s zero-shot setup for a fair comparison, consisting of a total of 3,000 samples considering three stance classes i.e., Favor, Against, and None. This includes 500 samples each from 6 different targets (see Appendix E). Additionally, we categorize the dataset into explicit and non-explicit samples. This categorization involves removing stop words and special characters from both texts and targets, lemmatizing all words, and subsequently comparing the text with the target. A sample is labeled as explicit if any target words are found in the corresponding text; otherwise, it is labeled as non-explicit. This process results in 1,804 explicit samples and 1,196 non-explicit samples (in Appendix E). It is important to note that because we compare our approach with TSE zero-shot, and they do not provided the code or sufficient details for reproducing the generation or selection of candidate targets, we are unable to use other stance detection datasets for benchmarking.
In our focused contribution, we utilize both proprietary and non-proprietary LLMs, including GPT-3.5 (Brown et al., 2020), Llama-3 (Touvron et al., 2023), and Mistral (Jiang et al., 2023), to address our two research questions (§1). [^3] Although we do not claim that the results or findings are true for all LLMs, experimenting with other proprietary LLMs is financially constraining for us. Model descriptions can be found in Appendix  B.
4.2 Evaluation Methods
To measure the generated target quality in OTSD task, we consider the following metrics:
BTSD: We fine-tune the BERTweet model (Nguyen et al., 2020) following the same setup as TSE (fine-tuning details in Appendix F). The SD score (F1-macro) is used as the evaluator for the generated targets. To assess the effectiveness of this metric, we experiment with different levels of target quality by modifying or removing words from the golden targets, substituting them with incorrect targets, and selecting random words from the vocabulary as targets. We find that the BERTweet F1 score consistently and significantly improved with higher-quality targets, in both explicit and non-explicit cases (Appendix C, Table 3). Therefore, the F1 score on stance detection serves as a reliable proxy for evaluating target generation quality in our dataset.
Human Evaluation: To evaluate the quality of each generated target and to further validate the BTSD metric, we conduct a small-scale human evaluation using 500 randomly selected samples (300 explicit and 200 non-explicit, with 16.67% from each stance label). Three independent domain experts are asked to assess the relevance of the generated targets by evaluating its relevance to the golden target in the tweet. Each annotator is provided with an individual file, ensuring that their evaluations remain independent of one another. The full agreement details and score distribution are included in Appendix I and J, respectively.
SemSim: In OTSD, since we do not rely on a predefined list of targets, the generated targets may not precisely match the gold targets but are semantically related. Therefore, we measure the semantic similarity between the generated and gold targets using BERT embeddings. Detailed steps can be found in Appendix G.
We use macro-average F1 for the task of SD.
TG+SD TG&SD
Model SemSim BTSD HEval SemSim BTSD HEval
Explicit
TSE-M 0.96 36.63 0.716 - - -
TSE-B 0.86 35.8 0.338 - - -
GPT-3.5 0.87 38.43 0.546 0.87 39.6 0.663
Llama-3 0.88 38.31 0.54 0.87 37.75 0.558
Mistral 0.85 38.47 0.586 0.87 39.02 0.65
Non-explicit
TSE-M 0.9 30.56 0.391 - - -
TSE-B 0.82 29.32 0.215 - - -
GPT-3.5 0.84 33.1 0.419 0.83 31.32 0.405
Llama-3 0.85 33.06 0.447 0.83 33.81 0.49
Mistral 0.83 30.18 0.386 0.83 34.12 0.495
Table 1: Comparison of TG performance between TSE and LLMs across various evaluation metrics. HEval represents human evaluation. TSE-M and TSE-B correspond to TSE-mapped and TSE-BestGen, respectively. The highest score for each metric across all models is highlighted in bold, while for BTSD and HEval, the better results between the two approaches are underlined.
5 Result Analysis
5.1 Target Generation
Following RQ#1, Table 1 shows that for the explicit case, Mistral achieves the highest BTSD score in the TG+SD setting, while GPT-3.5 leads in TG&SD. Overall, LLMs outperform TSE-BestGen in the BTSD metric and produce more relevant targets according to HEval scores. Although TSE-mapped has the highest SemSim and HEval scores, its BTSD score is lower due to generating either a perfect match or something entirely different (as illustrated in Figure 1). In contrast, the LLMs generate targets that are either highly related or partially related to the golden ones, with fewer irrelevant targets compared to TSE-mapped, leading to higher BTSD scores despite lower HEval. For non-explicit cases, LLMs outperform TSE-BestGen in both BTSD and HEval, with Llama-3 generating the most relevant targets (as indicated by the HEval score), even surpassing TSE-mapped. In the explicit case, TG&SD generally outperforms TG+SD across most LLMs, except for Llama, which performs better with TG+SD in BTSD. In non-explicit cases, GPT excels with TG+SD, while the other LLMs perform better with TG&SD. Among LLMs, GPT performs best overall in the explicit case, while Mistral leads in the non-explicit case. A manual inspection (on GPT-3.5’s 30 incorrect predictions) investigating the lower score in non-explicit cases reveals that the model struggles without sufficient context, often due to the way the TSE test set is constructed. This lack of context, such as implicit hints on the target or the context of other messages in the thread, causes the model to generate targets based solely on the provided text (example in Appendix H.1), leading to inaccuracies. Table 1 also confirms that higher HEval scores correlate with higher BTSD scores, validating BTSD as a reliable proxy for evaluating target quality.
Model TG+SD TG&SD
Explicit
TSE-mapped 38.1 -
TSE-BestGen 37.81 -
GPT-3.5 42.68 47.61
Llama-3 43.13 42.1
Mistral 46.15 46.32
Non-explicit
TSE-mapped 32 -
TSE-BestGen 31 -
GPT-3.5 32.66 33.94
Llama-3 33.36 31.9
Mistral 33.61 30.47
Table 2: Comparison of SD performance between TSE and LLMs, measured by F1-macro (in %). The highest scores across all models are highlighted in bold, with the best scores between the two approaches underlined.
5.2 Stance Detection
To address RQ#2, Table 2 shows that in stance detection, LLMs without a predefined list outperform TSE, which uses a predefined list, in both approaches (TG+SD and TG&SD) for explicit cases. In non-explicit cases, the scores are comparable in the TG+SD setting, but higher scores are observed for LLMs in the TG&SD setting. This is because, in explicit cases, LLMs generate better-quality targets than TSE-BestGen, leading to improved stance detection as evidenced in Table 1. For non-explicit cases, the holistic view provided by the TG&SD setting makes it easier for LLMs to infer stance from the text. Note that although LLM generated targets are better quality than TSE-BestGen, their stance labels can still be different from golden labels, resulting in low F1 scores. Our manual analysis revealed that these errors mainly arise from insufficient context in the input text, both explicit and non-explicit (see Appendix H.2 for examples). In both approaches, LLMs achieve much higher SD scores in explicit cases compared to non-explicit ones. Again, this is due to the incorrectly generated targets in the non-explicit case revealed by manual inspection. Incorrect targets correspond in a high percentage to incorrect stance labels. Most of targets were either completely incorrect or, in a few cases, contradictory to the real target, altering the stance.
6 Conclusion
LLMs outperform TSE in target generation for both explicit and non-explicit cases, despite TSE’s use of a predefined list. However, LLMs struggle more with generating targets when they are not explicitly mentioned in the text. LLMs also perform better in stance detection than TSE for both explicit and non-explicit cases. Additionally, LLMs show better stance detection performance when the target is explicitly mentioned compared to non-explicit cases. Finally, the single-step joint prompt approach (TG&SD) is more effective than the two-step approach (TG+SD) for both TG and SD in both explicit and non-explicit scenarios.
Limitation
We do not provide a specific coherence measure to evaluate the Open-Target Stance Detection task, as the target and the stance toward it are inherently interconnected. Due to resource constraints, our experiments were conducted on a small scale in terms of the models used. The human evaluation of target quality was also limited to 500 samples. Regarding potential data leakage during the pre-training of LLMs, it is possible that these models were exposed to some of our test data, leading to prior knowledge of certain entities. However, since the LLMs were not explicitly pre-trained on the same task, we continue to refer to our approach as zero-shot stance detection.
Appendix A Prompting Details
TG+SD (Two-Step Approach)
Prompt (TG):
You will be provided with a tweet, and your task is to generate a target for this tweet. A target should be the topic on which the tweet is talking. The target can be a single word or a phrase, but its maximum length MUST be 4 words. The output should only be the target, no other words. Do not provide any explanation but you MUST give an output, do not leave any output blank.
Prompt (SD):
Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a certain, specified target. Analyze the following tweet and determine its stance towards the provided target. If the stance is in favor of the target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE. Do not provide any explanation but you MUST give an output, do not leave any output blank. Only return the stance as a single word, and no other text.
Prompt Justification:
  • Clarity of Task: The task description explicitly states what the LLM needs to do—generate a target for a given tweet in TG prompt and for SD, it provides a clear and concise definition of stance classification. This ensures that the LLM understands the primary objective without ambiguity.
  • Focus on Relevance: By instructing the model that the target should be the topic of the tweet, it emphasizes the need for relevance, ensuring that the generated target accurately represents the tweet’s content.
  • Conciseness Requirement: Limiting the target to a maximum of 4 words ensures that the output remains concise and focused, which is crucial for the generated target to be compared fairly with the golden target.
  • Stance Options Provided: By specifying the possible stances (Favor, Against, None), the prompt sets clear categories for the model to choose from. This simplifies the models decision process and ensures the outputs are standardized.
  • Output Format Specification: Clearly specifying that the output should only be the target, with no additional words or explanations, helps to standardize the outputs, making them easier to evaluate.
  • Mandating Output Generation: By instructing the model not to leave any output blank and always to provide an answer, the prompt ensures robustness in responses, reducing the chance of missing data.
TG&SD (Single-Step Joint Approach)
Prompt (TG, SD):
Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a certain, specified target. Analyze the following tweet, generate the target for this tweet, and determine its stance towards the generated target. A target should be the topic on which the tweet is talking. The target can be a single word or a phrase, but its maximum length MUST be 4 words. If the stance is in favor of the target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE. If the stance is in favor of the generated target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE. The answer only has to be one of these three words: FAVOR, AGAINST, or NONE. Do not provide any explanation but you MUST give an output, do not leave any output blank. The output format should be: “‘Target: <target>, Stance: <stance>“‘.
Prompt Justification:
  • Integrated Task Definition: The prompt combines the definitions of both target generation and stance classification, ensuring the LLM understands the combined objective of generating a target and determining the stance towards it.
  • Clear and Concise Instructions: By clearly outlining the steps—first generating a target and then determining the stance towards it—the prompt guides the model through a logical sequence, reducing ambiguity and improving task performance.
  • Conciseness Requirement: Limiting the target to a maximum of 4 words ensures that the output remains concise and focused, which is crucial for the generated target to be compared fairly with the golden target.
  • Target Generation Constraints: Limiting the target to a maximum of 4 words ensures that the output remains concise and focused, which is crucial for the generated target to be compared fairly with the golden target.
  • Stance Classification Options: Providing explicit stance options (Favor, Against, None) simplifies the decision process for the model and ensures consistency in the outputs.
  • Mandatory Output Generation: By instructing the model to always provide an output and not leave any fields blank, the prompt ensures robustness and completeness in the responses, reducing the chance of missing data.
  • Output Format Specification: Specifying the output format as ’Target: <target>, Stance: <stance>’ standardizes the responses, making them easier to parse and evaluate.
  • Focus on Simplicity and Relevance: The prompt emphasizes that only the stance should be returned as a single word and no additional text, which maintains simplicity and relevance in the outputs.
Appendix B Model Descriptions
GPT-3.5-turbo-0125
is a variant of the GPT-3.5 series with 175 billion parameters which utilizes a transformer-based architecture. This model is decoder-only, similar to models like GPT-3 (Brown et al., 2020) and GPT-4 (Achiam et al., 2023), but with optimizations for efficiency and performance.
Llama-3-8B Instruct
is a decoder-only model in the LLaMA (Touvron et al., 2023) series with 8 billion parameters that follows the architectural principles of its predecessor, Llama-2 (Touvron et al., 2023), but has been instruction-tuned to improve its performance on a variety of tasks.
Mistral 7B Instruct
(Jiang et al., 2023) is a decoder-only model like Llama-2 (Touvron et al., 2023), but with different architectural features like Grouped-Query Attention and Sliding Window Attention. The Mistral model is 7 billion parameters and has been instruction tuned.
Appendix C BTSD Metric Justification
Input F1 (macro)
Explicit Non-explicit
Tweet only 31.02 30.37
-w/ Gold Target 41.67 34.65
-w/ altered GT 35.3 32.54
-w/ incorrect Target 25.58 23.58
-w/ random vocab 18.66 14
Table 3: Macro avg. F1-Score of BertTweet on Stance Detection with different quality levels of the input target.
Table 3 contains the results of the stratified experiments with different measure stance detection F1 score when giving as input the text and a target, varying the quality of this latter. The numbers are significantly different.
Appendix D Hardware Setting
All of the experiments were run on a computer with Ubuntu 22.04 Linux with x64 CPU with 4 cores, 32 GB of RAM and one NVIDIA V100 GPU.
Appendix E Dataset Samples
Targets: Creationism (Somasundaran and Wiebe, 2010), Gay Rights (Somasundaran and Wiebe, 2010), Climate Change is a Concern (Mohammad et al., 2016), MeToo Movement (Gautam et al., 2020), Merger of Disney and Fox (Conforti et al., 2020), and Lockdown in New York State (Miao et al., 2020). [NEED TO EDIT]
Appendix F BTSD Fine-tuning
Following TSE, we consider 4 datasets combined i.e. SemEval (Mohammad et al., 2016), AM (Stab et al., 2018), COVID-19 (Glandt et al., 2021), P-Stance (Li & Caragea, 2021), containing 19 targets, splitting samples (70-80%, 10-15%, 15-20%) depending on the dataset size into train, dev, and test respectively, maintaining the distribution from each of the targets. We will clarify it further in our revised version. [NEED TO EDIT]
Appendix G SemSim Details
Specifically, we employ the bert-base-uncased (Devlin et al., 2018) model to obtain contextual encoded representations of the target words. After performing a mean pooling operation on these representations, we apply cosine similarity (Singhal, 2001) to the encoded sequences to obtain the similarity score.
Appendix H Explanation with Samples in TG and SD
H.1 Target Generation
For instance, in the post "I mean, this reads to me like anyone who’s an alcoholic doesn’t go to heaven…", GPT-3.5 generated "Sinful behavior and heaven’s eligibility" as the target instead of the gold target "gay rights".
H.2 Stance Detection
For instance, in the explicit sample text, "Q. from audience: Malcom X said the most disrespected person in America is the black woman, has #MeToo changed this? Gender has been constructed in the image of the white woman, the idea of a level playing field, white women are often in better positions, says @HeidiMirza #CHevents", GPT-3.5 generated the target "Gender roles", which, while related, fails to capture the specific stance-related target "MeToo movement". This mistake reflects the model’s need for broader context to establish the correct link. Similarly, in the non-explicit sample "@DrOakley1689 @PhxApologetics Maybe they just misused their resources and their societies collapsed #deforestation #SemST", GPT-3.5 generated "Environmental degradation" as the target, while the gold target was "Climate change is a concern". Here, the generated target, though relevant as a subtopic, misses the broader context necessary to align with the gold target.
Appendix I Human Annotator Agreement
Dataset Description: The file includes a Tweet column with various Twitter posts. The Gold Target column represents the true target, indicating the primary topic or issue the tweet addresses, which in turn informs the stance noted in the Gold Stance column. For each tweet, there are eight generated targets (in columns T1 through T8) produced by different models, whose names remain anonymous for unbiased annotation. Additionally, there are empty cells (e.g., in the T1 Score column) designated for storing human evaluation scores corresponding to each generated target.
Definition of a Relevant Target: A generated target in the context of the target generation task is considered relevant if it accurately captures or closely aligns with the main topic, issue, or entity identified by the golden target in a given tweet. It should reflect the key subject matter or concern that the golden target represents, maintaining coherence with the overall context and meaning of the tweet.
Task Description: Your task is to assess the quality of each generated target by evaluating its relevance to the golden target in the tweet. You will do this by answering the following question and assigning a score to each generated target. Scoring criteria and examples are provided below.
“How closely does the generated target relate to the golden target in the tweet?”
Scoring Scale: 0 to 1 (Low to High)
  • 0 – Not Related: The generated target is entirely unrelated to the golden target. It does not reference or express anything connected to the intended topic.
  • 0.5 – Partially or Indirectly Related: The generated target has some relevance to the golden target but does not directly match the main concept. It may address a broader or narrower aspect of the topic, such as a parent topic, a subtopic, or a tangentially connected idea that is not fully aligned with the golden target’s core meaning.
  • 1 – Completely Related: The generated target is highly relevant to the golden target. It either uses a synonymous term, is semantically similar to the golden one, or conveys the same underlying idea, topic, or issue as the golden target. This means that the generated target captures the essence or main concept of the golden target, even if it uses different wording. For example, if the golden target is "climate change," a conceptually similar term could be "global warming" or "environmental crisis," as both refer to the broader issue of environmental concerns related to climate.
Example Scoring:
  1. 1.
    Tweet: Attempts to conceal the creationism-evolution controversy from students are dogmatic promotions of evolution. Not since blasphemy laws has competitive expression of thought been illegalized, and this is what evolutionists want to accomplish. This is evidenced by none other than the title of an evolutionist argument on this very page: "Schools should not teach theories that are completely at odds with each other",
    Gold Target: creationism,
    T1: creationism-evolution controversy, T1 Score: 1,
    T2: Education controversy, T2 Score: 0.5.
  2. 2.
    Tweet: If, on a supernatural level, lust of any sort counts the same consummated or unconsummated, why not just go ahead and consummate. You can’t get in any worse trouble, and if it makes you happy…,
    Gold Target: gay rights,
    T1: consummation, T1 Score: 0.5,
    T2: superstition, T2 Score: 0.
Appendix J Human Annotation Score Distribution
Figure 1: Distribution of human evaluation scores for target relevance assessment (0 – Not Related, 0.5 – Partially or Indirectly Related, 1 – Completely Related). The final score for each sample is determined by the majority vote among the three annotators. In cases of a tie, the average score of 0.5 is assigned. Complete details on human annotator agreement are provided in Appendix I.